OpenStudy (anonymous):

I need help writing an equation in point-slope form for the line through the given point with the given slope (-4, -2); m=-1/2 I think the answer is either y-2 = -1/2 (x-4) or y+2 = -1/2(x-4) ??

OpenStudy (freckles):

point slope form is y-y1=m(x-x1) where (x1,y1) is the point given and m is the slope

OpenStudy (freckles):

you are given m is -1/2 so you did good there

OpenStudy (freckles):

you are given x1 is -4 and y1 is -2

OpenStudy (freckles):

so if you put those in then you have y-(-2)=-1/2 (x-(-4)) correct?

OpenStudy (freckles):

but you can simplify that a little

OpenStudy (freckles):

since -*-=+

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

That is where I was confused on if it was minus 2 or +2

OpenStudy (freckles):

but you understand now that if you have a -*- you have a +

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (freckles):

ok one more

OpenStudy (anonymous):

(5, -6); m=-0.38 I think it is y+6=-0.38(x+5)

OpenStudy (freckles):

ok well again point slope form is y-y1=m(x-x1) everything is right except for your x1

OpenStudy (freckles):

x1=5

OpenStudy (freckles):

not -5

OpenStudy (anonymous):

wait didn't I say y + 6 = -0.38(x+5)?

OpenStudy (freckles):

yes but you said x was 5

OpenStudy (freckles):

but you entered in -5

OpenStudy (freckles):

\[y-y_1=m(x-x_1) \\ (x_1,y_1)=(5,-6) \text{ and } m=-0.38 \\ y-(-6)=-0.38(x-5)\]

OpenStudy (anonymous):

Okay! Thanks so very much!

OpenStudy (freckles):

examples: \[(1,1) ; y-1=m(x-1) \\ (1,-1);y-(-1)=m(x-1) \\ (-1,1); y-1=m(x-(-1)) \\ (-1,-1); y-(-1)=m(x-(-1)) \\ \text{ simplifying } \\ (1,1); y-1=m(x-1) \\ (1,-1); y+1=m(x-1) \\ (-1,1); y-1=m(x+1) \\ (-1,-1); y+1=m(x+1)\]

OpenStudy (freckles):

I guess if it really confuses you think opposite when plugging into the point slope form so if x=1 then I will x-1 and if i have x=-1 then I will have x+1 and if I have y=1 then I will have y-1 and if I have y=-1 then I will have y+1 but the only thing about this is if you memorize this it might also make other things more confusing in the future

OpenStudy (freckles):

so if I have the point (7,-4) with slope -2 then I have y+4=-2(x-7)

OpenStudy (anonymous):

Okay, but that did help. I wont memorize it! I don't like math, it is very much like a foreign language!!!

OpenStudy (freckles):

i bet one day you will be all giggly to do math

OpenStudy (anonymous):

lol I doubt it! :))

OpenStudy (anonymous):

Thank You for your help! Have a Merry Christmas! :)

OpenStudy (freckles):

Happy Holidays to you as well.